OpenStudy (anonymous):

whats the cosine & sine of 30 ?

OpenStudy (anonymous):

sine30 = 1/2 ....cos 30 =(sqrt(3))/2

OpenStudy (anonymous):

wolframalpha.com great for studying!

Parth (parthkohli):